# The sup-norm problem for newforms of large level on $\operatorname{PGL}(n)$

### v1

@@ 6.4. Counting at different cusps
By Theorem III.12 in [New72], there is $h\in\operatorname{SL}_{n}(\mathbb{Z})$ such that
$h\gamma h^{-1}=\begin{pmatrix}m&\ast&\ast\\ &\ddots&\ast\\ &&m\end{pmatrix}$ (6.4)
is upper triangular with $m$ on the diagonal. Indeed, $\chi_{\gamma}$ splits into linear factors and thus the blocks in [New72, Thm. III.12] are one dimensional.
@@
recalling the definition of $\beta(z)$ in Section 5.2 and that $h\in\Gamma_{0}(N)$. By assumption, $z\in\mathcal{L}^{\prime}(B,A)$, which by Table 1 implies that $\beta(z)=N^{1/n}\lambda_{1}$ for $\lambda_{1}$ the first successive minimum of $L_{z^{\prime}}$. Lemma 6.1 shows now that $\beta(z)\asymp N^{1/n}$. As such, we have
@@
This is now a similar situation in the proofs of the counting results Proposition 6.2 and Proposition 6.4, except that $d$ might be large. However, the last row of $\eta$ is already fixed to be $m\cdot e_{n}$. As in Proposition 6.4, the assumption $L\ll N^{1/n^{2}-\varepsilon}$ and the bound above on the entries of $a(y)$ imply that the last row of $\eta$ determines the whole matrix. Therefore, $\eta=m\cdot\operatorname{id}_{n}$ and so, undoing conjugation, $\gamma=m\cdot\operatorname{id}_{n}$. The case $\gamma\in H(z^{\prime},m^{n},N)$ follows analogously. What changes is, for instance, that $\beta(z^{\prime})=N^{1/n}\lambda_{1}$ for $\lambda_{1}$ the minimum of $L_{z}$. We then continue by using Lemma 6.1 again. ∎
@@
$\overline{W}_{n}\cong\Gamma_{0}(N)\cap W_{n}\backslash W_{n}.$
and note also that $|\overline{W}_{n}|=n$.
@@
Let $\xi\in\operatorname{SL}_{n}(\mathbb{Z})$ and let $(a_{1},\ldots,a_{n})$ be the first column of $\xi$, a primitive vector in $\mathbb{Z}^{n}$. First, we reduce $a_{n}$ to either 00 or $1$ by acting from the left by $\Gamma_{0}(N)$. Indeed, assume that $\gcd(a_{n},N)=1$. Then the vector $(Na_{1},\ldots,Na_{n-1},a_{n})$ is also primitive. Therefore there is a primitive $(b_{1},\ldots,b_{n})\in\mathbb{Z}^{n}$ such that
@@
From this it is clear that $\gcd(N,b_{n})=1$ so that $(Nb_{1},\ldots,Nb_{n-1},b_{n})$ is primitive. Let $\gamma\in\operatorname{SL}_{n}(\mathbb{Z})$ be a matrix with the latter as its last row. Then $\gamma\in\Gamma_{0}(N)$ and $\gamma\xi$ has last row of the form $(1,\ast,\ldots,\ast)$. Since $N$ is prime, negating the assumption above means that $N\mid a_{n}$. Now let $d=\gcd(a_{1},\ldots,a_{n-1})$. Then $\gcd(a_{n},d)=1$ and there exists a primitive vector $(b_{1},\ldots,b_{n-1})$ such that
@@
The vector $(a_{n}b_{1},\ldots,a_{n}b_{n-1},-d)$ is primitive by the observations above, so there is $\gamma\in\operatorname{SL}_{n}(\mathbb{Z})$ with this vector as its last row. Again, $\gamma\in\Gamma_{0}(N)$ since $N\mid a_{n}$ and the last row of $\gamma\xi$ has the form $(0,\ast,\ldots,\ast)$. Assume now that $\xi$ has last row of the form $(1,\ast,\ldots,\ast)$. It is clear that we can multiply $\xi$ from the right by a matrix in $U_{n}(\mathbb{Z})$ such that the resulting last row is simply $(1,0,\ldots,0)$. Call this new matrix $\xi$ again and take $w\in W_{n}$ a permutation matrix with the same last row (for instance the so-called long Weyl element). In other words, $e_{n}\xi=e_{n}w$, where $e_{n}$ is the $n$-th standard basis vector $(0,\ldots,0,1)$. The matrix $w\xi^{-1}$ preserves $e_{n}$ so it must have $e_{n}$ as its last row. In particular, $w\xi^{-1}\in\Gamma_{0}(N)$ and we are done in this case. On the other hand, let $\xi$ have last row of the form $(0,\ast,\ldots,\ast)$. Using the embedding of $\operatorname{SL}_{n-1}(\mathbb{Z})$ in the upper left corner of $\Gamma_{0}(N)$, we may modify $\xi$ so that its first column is of the form $(1,0,\ldots,0)$, by similar arguments. This now allows an inductive procedure, considering the lower right $n-1\times n-1$ block of $\xi$. We see that one can always reduce the last row of $\xi$ to be a standard basis vector and the paragraph above shows how to obtain a permutation matrix from $\xi$. To check that no two such representatives in $\overline{W}_{n}$ produce the same double coset is easy. For $w_{1},w_{2}\in W_{n}$, if $w_{1}=\gamma w_{2}u$ with $\gamma\in\Gamma_{0}(N)$ and $u\in U_{n}(\mathbb{Z})$, then $\gamma=w_{1}u^{-1}w_{2}$. One now computes the shape of $U_{n}(\mathbb{Z})$ transformed by permutation of rows and of columns. We leave out the details of this argument. ∎
@@
[…]
$\Delta_{n-1}(\gamma)=q^{(n-1)(n-2)}$
@@
Let $m=pq^{n-1}$. As in (6.4), there is $h\in\operatorname{SL}_{n}(\mathbb{Z})$ such that $h\gamma h^{-1}$ is upper-triangular with diagonal $(m,\ldots,m)$. By Lemma 6.8 we can write $h^{-1}=\sigma^{-1}wu^{-1}$ with $\sigma\in\Gamma_{0}(N)$, $u\in U_{n}(\mathbb{Z})$, and $w\in\overline{W}_{n}$. Next, conjugating by $u$, we easily see that
$w^{T}\sigma\gamma\sigma^{-1}w=\begin{pmatrix}m&\ast&\ast\\ &\ddots&\ast\\ &&m\end{pmatrix}=:\eta$ (6.5)
is also of the same form. Now if $w=w_{n}=\operatorname{id}_{n}$, meaning that the last row of $w$ is $e_{n}$ as in Remark 6.9, we are done by Lemma 6.7. The latter implies that $\gamma=pq^{n-1}\cdot\operatorname{id}_{n}$, which does not have the required determinantal divisors and leads to a contradiction. If $w=w_{1}$ is the long Weyl element, we apply the Fricke involution. By transposing the condition
$z^{-1}\gamma z=O(m).$
@@
lies in $H(\tilde{\sigma}z^{\prime},m^{n},N)$ with some $\tilde{\sigma}\in\Gamma_{0}(N)$. Next, observe that
@@
and thus $\gamma=m\cdot\operatorname{id}_{n}$, which is a contradiction again. Finally, let $w=w_{k}$ with $1<k<n$. Notice first that (6.5) and the congruences modulo $N$ satisfied by $\gamma$ and $\sigma$ imply that the $k$-th row of $\eta$ also satisfies congruences. Indeed, $w_{k}\eta w_{k}^{T}$ is a matrix of the $\Gamma_{0}(N)$ shape. More precisely, $N\mid\eta_{kj}$ for $j>k$. Since $k<n$, we have in particular $N\mid\eta_{k,k+1}$. Let us now assume that the superdiagonal of $\eta$ only contains non-zero elements. That is, $\eta_{j,j+1}\neq 0$ for all $1\geq j<n$. Recall the condition
@@
We can rewrite $w^{T}\sigma z=n\cdot a$ in Iwasawa coordinates (indeed, conjugating by an orthogonal matrix leaves $O(m)$ invariant), denoting the $y$-coordinates as usual. It is now a common and important observation that the superdiagonal of upper triangular matrices enjoys a certain additive abelian-like property with respect to matrix multiplication. This observation or direct computation should convince the reader that
@@
$y_{1}\cdots y_{n-1}\gg\frac{N}{m^{n-1}}\gg N^{1/n+\varepsilon}$
using the assumption $L\ll N^{1/n^{2}-\varepsilon}$ and that $m\asymp L^{n}$. We return now to a technique used in the proof of Lemma 6.7. We observe again that
@@
$\left\lVert e_{2}\wedge\cdots\wedge e_{n}\right\rVert_{w^{T}\sigma z}=\left\lVert e_{2}\wedge\cdots\wedge e_{n}\right\rVert_{\sigma z}\geq\beta(z)\asymp N^{1/n}.$
Therefore, as in the proof of the aforementioned lemma, we obtain that
$dy_{1}\cdots y_{n-1}\ll N^{1/n}.$
Recall also that $d=\left\lVert e_{n}\right\rVert_{w^{T}\sigma z}$, and since $z$ defines a balanced lattice, $d\gg 1$. Therefore
@@
which constitutes a contradiction to the previous paragraph for large enough $N$. We deduce that the superdiagonal of $\eta$ must contain some zero. It is now straight-forward to prove that $m$ divides $\Delta_{n-1}(\eta)$. Indeed, the only $(n-1)\times(n-1)$ minor that is not obviously divisible by $m$ is the upper right minor, formed by removing the first column and the last row of $\eta$. Proving the claim here is an easy exercise in Laplace, or cofactor, expansion. Observe now that the invariance properties of determinantal divisors (see [New72, Thm. II.8]) imply that
@@
Notice that the case $n=2$ does not involve any intermediate Weyl elements. Indeed, there are only two cusps and both reduce as above to counting upper-triangular matrices directly. A more general result (for square-free levels) is contained in a slightly different language in [HT12, Lemma 4.1].
@@
Let $n$ and $N$ be a prime, and let $z\in\mathbb{H}$ reduce to a compactum $\Omega$. Assume that $z$ has Iwasawa coordinates
[…]
for $i=2,\ldots,n-1$, and $d=\alpha(z)=\left\lVert e_{n}\right\rVert_{z}$, and satisfies
$z\in\mathcal{L}(B,A)\cap\mathcal{L}^{\prime}(B,A).$
Let $L\ll N^{1/n^{2}-\varepsilon}$ and assume that $N\gg_{\Omega,\varepsilon}1$ is large enough. Then
@@ 7. Final steps
7.1. The Fourier bound
To prove a bound in the first domain given by the reduction in Proposition 5.2, we use the Whittaker expansion and bounds for the first Fourier coefficient of newforms of level $N$. Here we state a version of the bound that is unconditional, yet sufficient for our purposes.

### v2

@@ 6.4. Counting at different cusps
By Theorem III.12 in [35], there is $h\in\operatorname{SL}_{n}(\mathbb{Z})$ such that
(6.4) $h\gamma h^{-1}=\begin{pmatrix}m&\ast&\ast\\ &\ddots&\ast\\ &&m\end{pmatrix}$
is upper triangular with $m$ on the diagonal. Indeed, $\chi_{\gamma}$ splits into linear factors and thus the blocks in [35, Thm. III.12] are one dimensional.
@@
recalling Remark 2.4, the definition of $\beta(z)$ in Section 5.3, and that $h\in\Gamma_{0}(N)$. We are assuming that $z$ satisfies the conditions in case (2.b) of Proposition 5.3. Thus, $z\in\mathcal{L}^{\prime}(B,A)$, which by Table 1 implies that $\beta(z)=N^{1/n}\lambda_{1}(L_{z^{\prime}})$. The assumptions also give that $\lambda_{1}(L_{z^{\prime}})\gg N^{-1/2n}$ and so $\beta(z)\gg N^{1/2n}$. As such, we have
$dy_{1}\ll dy_{1}y_{2}\ll\ldots\ll dy_{1}\cdots y_{n-1}\ll N^{-1/2n}.$
We are now again in a similar situation as in the counting results Proposition 6.2 and Proposition 6.4. However, the last row of $\eta$ is already fixed to be $m\cdot e_{n}$. As in Proposition 6.4, the assumption $L\ll N^{1/2n^{2}-\varepsilon}$, the bound above on the entries of $a(y)$, and the assumption that $L_{z}$ is balanced imply that the last row of $\eta$ determines the whole matrix. Therefore, $\eta=m\cdot\operatorname{id}_{n}$ and so, undoing conjugation, $\gamma=m\cdot\operatorname{id}_{n}$.
The case $\gamma\in H(z^{\prime},m^{n},N)$ follows analogously. Here $z\in\mathcal{L}(B,A)$, but $L_{z}$ is also balanced, so $\beta(z^{\prime})=N^{1/n}\lambda_{1}(L_{z})\asymp N^{1/n}$. Therefore, the $y$-coordinates of $hz^{\prime}$ satisfy the stronger bound
@@
On the other hand, $L_{z^{\prime}}$ is potentially unbalanced, yet $\lambda_{1}(L_{z^{\prime}})\gg N^{-1/2n}$. By following the proof of Proposition 6.4, we are considering $z^{\prime}$-lattice points inside balls of radius $L^{n}\cdot dy_{1}\cdots y_{k}$ for some $k$. The assumption on $L$ bounds the radius by $N^{1/2n-\varepsilon-1/n}=N^{-1/2n-\varepsilon}$, which is smaller than the shortest length in $L_{z^{\prime}}$ for $N$ large enough. Thus, there is at most on lattice point in such a ball, and the proof continues as usual. ∎
@@
[…]
Note also that $|\overline{W}_{n}|=n$.
@@
Let $\xi\in\operatorname{SL}_{n}(\mathbb{Z})$ and let $(a_{1},\ldots,a_{n})$ be the first column of $\xi$, a primitive vector in $\mathbb{Z}^{n}$. First, we reduce $a_{n}$ to either 0 or $1$ by acting from the left by $\Gamma_{0}(N)$.
Indeed, assume that $\gcd(a_{n},N)=1$. Then the vector $(Na_{1},\ldots,Na_{n-1},a_{n})$ is also primitive. Therefore, there is a primitive $(b_{1},\ldots,b_{n})\in\mathbb{Z}^{n}$ such that
@@
From this it is clear that $\gcd(N,b_{n})=1$ so that $(Nb_{1},\ldots,Nb_{n-1},b_{n})$ is primitive. Let $\gamma\in\operatorname{SL}_{n}(\mathbb{Z})$ be a matrix with the latter as its last row. Then $\gamma\in\Gamma_{0}(N)$ and $\gamma\xi$ has last row of the form $(1,\ast,\ldots,\ast)$.
Since $N$ is prime, negating the assumption above means that $N\mid a_{n}$. Now let $d=\gcd(a_{1},\ldots,a_{n-1})$. Then $\gcd(a_{n},d)=1$ and there exists a primitive vector $(b_{1},\ldots,b_{n-1})$ such that
@@
The vector $(a_{n}b_{1},\ldots,a_{n}b_{n-1},-d)$ is primitive by the observations above, so there is $\gamma\in\operatorname{SL}_{n}(\mathbb{Z})$ with this vector as its last row. Again, $\gamma\in\Gamma_{0}(N)$ since $N\mid a_{n}$ and the last row of $\gamma\xi$ has the form $(0,\ast,\ldots,\ast)$.
Assume now that $\xi$ has last row of the form $(1,\ast,\ldots,\ast)$. It is clear that we can multiply $\xi$ from the right by a matrix in $U_{n}(\mathbb{Z})$ such that the resulting last row is simply $(1,0,\ldots,0)$. Call this new matrix $\xi$ again and take $w\in W_{n}$ a permutation matrix with the same last row (for instance the so-called long Weyl element). In other words, $e_{n}\xi=e_{n}w$, where $e_{n}$ is the $n$-th standard basis vector $(0,\ldots,0,1)$. The matrix $w\xi^{-1}$ preserves $e_{n}$ so it must have $e_{n}$ as its last row. In particular, $w\xi^{-1}\in\Gamma_{0}(N)$ and we are done in this case.
On the other hand, let $\xi$ have last row of the form $(0,\ast,\ldots,\ast)$. Using the embedding of $\operatorname{SL}_{n-1}(\mathbb{Z})$ in the upper left corner of $\Gamma_{0}(N)$, we may modify $\xi$ so that its first column is of the form $(1,0,\ldots,0)$, by similar arguments. This now allows an inductive procedure, considering the lower right $(n-1)\times(n-1)$ block of $\xi$. We see that one can always reduce the last row of $\xi$ to be a standard basis vector and the paragraph above shows how to obtain a permutation matrix from $\xi$.
To check that no two such representatives in $\overline{W}_{n}$ produce the same double coset is easy. For $w_{1},w_{2}\in W_{n}$, if $w_{1}=\gamma w_{2}u$ with $\gamma\in\Gamma_{0}(N)$ and $u\in U_{n}(\mathbb{Z})$, then $\gamma=w_{1}u^{-1}w_{2}$. One now computes the shape of $U_{n}(\mathbb{Z})$ transformed by permutation of rows and of columns. We leave out the details of this argument. ∎
@@
Assume the same conditions as in Corollary 6.5. For $N$ large enough, the set of matrices $\gamma$ possibly occurring in $H_{\text{par}}(z,(pq^{n-1})^{n},N)$ for some primes $\asymp L, pqp\neq q$, satisfying additionally that
$\Delta_{n-1}(\gamma)=q^{n(n-2)}$
@@
Let $m=pq^{n-1}$. As in (6.4), there is $h\in\operatorname{SL}_{n}(\mathbb{Z})$ such that $h\gamma h^{-1}$ is upper-triangular with diagonal $(m,\ldots,m)$. By Lemma 6.8 we can write $h^{-1}=\sigma^{-1}wu^{-1}$ with $\sigma\in\Gamma_{0}(N)$, $u\in U_{n}(\mathbb{Z})$, and $w\in\overline{W}_{n}$.
Next, conjugating by $u$, we easily see that
(6.5) $w^{T}\sigma\gamma\sigma^{-1}w=\begin{pmatrix}m&\ast&\ast\\ &\ddots&\ast\\ &&m\end{pmatrix}=:\eta$
is also of the same form. Now if $w=w_{n}=\operatorname{id}_{n}$, meaning that the last row of $w$ is $e_{n}$ as in Remark 6.9, we are done by Lemma 6.7. The latter implies that $\gamma=pq^{n-1}\cdot\operatorname{id}_{n}$, which does not have the required determinantal divisors and leads to a contradiction.
If $w=w_{1}$ is the long Weyl element, we apply the Fricke involution. By transposing the condition
$z^{-1}\gamma z=O(m),$
@@
lies in $H(\tilde{\sigma}z^{\prime},m^{n},N)$ with some $\tilde{\sigma}\in\Gamma_{0}(N)$.
Next, observe that
@@
and thus $\gamma=m\cdot\operatorname{id}_{n}$, which is a contradiction again.
Finally, let $w=w_{k}$ with $1<k<n$. Notice first that (6.5) and the congruences modulo $N$ satisfied by $\gamma$ and $\sigma$ imply that the $k$-th row of $\eta$ also satisfies congruences. Indeed, $w_{k}\eta w_{k}^{T}$ is a matrix of the $\Gamma_{0}(N)$ shape. More precisely, $N\mid\eta_{kj}$ for $j>k$. Since $k<n$, we have in particular $N\mid\eta_{k,k+1}$.
Let us now assume that the superdiagonal of $\eta$ only contains non-zero elements. That is, $\eta_{j,j+1}\neq 0$ for all $1\leq j<n$. Recall the condition
@@
We can rewrite $w^{T}\sigma z=n(x)\cdot a(y)$ in Iwasawa coordinates (indeed, conjugating by an orthogonal matrix leaves $O(m)$ invariant), denoting the $y$-coordinates as usual. It is now a common and important observation that the superdiagonal of upper triangular matrices enjoys a certain additive abelian-like property with respect to matrix multiplication. This observation or direct computation should convince the reader that
@@
$y_{1}\cdots y_{n-1}\gg\frac{N}{m^{n-1}}\gg N^{1/2}$
using the assumption $L\ll N^{1/2n^{2}-\varepsilon}$ and that $m\asymp L^{n}$.
We return now to a technique used in the proof of Lemma 6.7. We observe again that
@@
$\left\lVert e_{2}\wedge\cdots\wedge e_{n}\right\rVert_{w^{T}\sigma z}=\left\lVert e_{2}\wedge\cdots\wedge e_{n}\right\rVert_{\sigma z}\geq\beta(z).$
Recall that, in case (2.b) of Proposition 5.2, we have $\beta(z)=N^{1/n}\lambda_{1}(L_{z^{\prime}})$. The assumptions also give that $\lambda_{1}(L_{z^{\prime}})\gg N^{-1/2n}$, so $\beta(z)\gg N^{1/2n}$. We obtain that
$dy_{1}\cdots y_{n-1}\ll N^{-1/2n}.$
Recall also that $d=\left\lVert e_{n}\right\rVert_{w^{T}\sigma z}$, and since $z$ defines a balanced lattice, $d\gg 1$. Therefore,
@@
which constitutes a contradiction to the previous paragraph for large enough $N$.
We deduce that the superdiagonal of $\eta$ must contain some zero. It is now straight-forward to prove that $m$ divides $\Delta_{n-1}(\eta)$. Indeed, the only $(n-1)\times(n-1)$ minor that is not obviously divisible by $m$ is the upper right minor, formed by removing the first column and the last row of $\eta$. Proving the claim here is an easy exercise in Laplace, or cofactor, expansion.
Observe now that the invariance properties of determinantal divisors (see [35, Thm. II.8]) imply that
@@
Notice that the case $n=2$ does not involve any intermediate Weyl elements. Indeed, there are only two cusps and both reduce as above to counting upper-triangular matrices directly. A more general result (for square-free levels) is contained in a slightly different language in [21, Lemma 4.1].
@@
Let $n$ and $N$ be a prime, and let $z\in\mathbb{H}$ satisfy the conditions of case (2.b) in Proposition 5.2. Let $L\ll N^{1/2n^{2}-\varepsilon}$ and assume that $N\gg_{\Omega,\varepsilon}1$ is large enough. Then
@@ 7. Final steps
First, assume Hypothesis (1.3). Proposition 3.3 and Proposition 6.2 together with the prime number theorem imply that
under the assumptions on $z$ specified in Proposition 6.2. Optimising the size of $L$, we choose $L=N^{1/2n}$. In this case, we have
$\phi(z)\ll L^{-1/4+\varepsilon}\ll N^{-1/8n+\varepsilon}.$
These bounds are valid on the subsets of $\mathbb{H}$ corresponding to case (2.b) in Proposition 5.2. As remarked at the beginning of Section 5, by applying the Fricke involution, these now extend to a subset of $\Omega_{N}^{\natural}$ that we call $\Omega_{N}$, the bulk. The complement, $C_{N}$, is given by the points corresponding to the cases (1) and (2.a). These form a set with $y$-coordinates as in (1.1), thus away from the bulk of $\Omega_{N}^{\natural}$, of volume bounded from above by $N^{(n-1)(1-1/2n^{4})}$. We use here that the volume form is invariant under the Fricke involution. One can see this, for instance, from the invariance of the Riemannian metric under $g\mapsto g^{-T}$.
for $L\ll N^{1/2n^{2}-\varepsilon}$. Maximising $L$, we get
$\phi(z)\ll L^{-1/2+\varepsilon}\ll N^{-1/4n^{2}+\varepsilon}.$
8. Appendix: The Fourier bound
Consider the cases (1) and (2.a) in Proposition 5.2. They give lower bounds for $y$-coordinates, away from the bulk $y_{1}\asymp 1/N$ and $y_{i}\asymp 1$ for $i\geq 2$. In an ideal situation, as in the case of $\operatorname{PGL}(2)$, the Fourier bound would be effective already in such regions. Indeed, as in [21, Lemma 4], there we have $\phi(x+iy)\ll N^{-1/2+\varepsilon}y^{-1/2}$, and we obtain a sub-baseline bound as soon as $y\geq N^{-1+\delta}$. The limit of this tool is precisely the bulk $y\asymp 1/N$.
However, the state of the art does not provide such useful Fourier bounds. This is due to a lack of control in the Whittaker expansion, particularly for the dual form of $\phi$, as well as the weakness of bounds on residues of Rankin-Selberg $L$-functions in higher rank.
For a demonstration of the current technology, we prove a bound in case (1). It is unconditional, yet even under a no-Siegel-zero assumption there are serious obstacles to generalising and improving it. This goes beyond the scope of this paper.